OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -6, an = 4 • an-1

jimthompson5910 (jim_thompson5910):

The first term is a1 = -6. This is given The next term must be found using the rule an = 4 • a(n-1)

jimthompson5910 (jim_thompson5910):

This rule says "multiply the previous term by 4 to get the next term" So a2 = 4*a1 ----> a2 = 4*(-6) = -24

jimthompson5910 (jim_thompson5910):

What is a3?

OpenStudy (jdoe0001):

\(\Large \bf a_1=-6\qquad a_{\color{red}{ n}}=4\cdot a_{{\color{red}{ n}}-1}\\ \quad \\ a_{\color{red}{ 2}}=4\cdot a_{{\color{red}{ 2}}-1}\qquad a_{\color{red}{ 3}}=4\cdot a_{{\color{red}{ 3}}-1}\qquad ...\)

OpenStudy (anonymous):

WHAT